OpenStudy (clamin):

Please help!! Medal!! Find the 500th term of the sequence below. 24,30,36,42,48,.... a.3018 b.3024 c.3042 d.2994

OpenStudy (anonymous):

the first term is 24.the next term is 24+6. so common difference is 6. then use this formula to find the 500th term.\[a_{n}=a _{1}+(n-1)d\]

OpenStudy (clamin):

an=a1 +(n-1) 6

OpenStudy (clamin):

B

OpenStudy (anonymous):

an=24+499*6

OpenStudy (clamin):

How did u get 499?? @amirreza1870

OpenStudy (anonymous):

your n is 500 so n-1 is 499.

OpenStudy (clamin):

Does it always to be subtract 1? @amirreza1870

OpenStudy (anonymous):

ye.s if you want to find 150th (n) term of your arithmetic sequence you should multiply common difference by 149 (n-1) and add the value of this multiplication to the first term. an=a1+(n-1)d